# Chronic Kidney Disease Prognosis Prediction Using Transformer

### v1

@@ 1 Introduction
Chronic kidney disease (CKD) is a progressive condition affecting millions worldwide, up to 10% of the global population, often leading to end-stage renal failure and significant morbidity [1]. Accurate prognosis prediction is critical for guiding treatment strategies, optimizing resource allocation, and improving patient outcomes. Traditional prognostic models rely on statistical approaches or shallow machine learning techniques, which often struggle to capture the complex temporal and contextual relationships inherent in longitudinal clinical data.
Recent advances in deep learning, particularly transformer architectures[2], have revolutionized sequence modeling by leveraging self-attention mechanisms to capture long-range dependencies and contextual patterns. Unlike recurrent models like Recurrent Neural Network (RNN), transformers enable parallel processing and dynamic weighting of features, making them well-suited for heterogeneous healthcare data such as electronic health records, laboratory results, and demographic information.
Previous studies on prognosis prediction of chronic kidney disease using deep learning have primarily employed recurrent neural networks (RNNs) [3] and transformer-based architectures [4][5][6]. However, these approaches have not fully explored the integration of transformer models with laboratory values. Recently, foundational models have attempted to incorporate laboratory data, as demonstrated by Ethos [7], which discretized continuous lab values into 10 quantiles for model input. Building on this idea, we adapted the decoder-based architecture of Ethos into an encoder-based model, achieving significant performance improvements in prognosis prediction.
@@ 2.3 Patient Sequences and Tokenization
As depicted in Figure 2, the core of the method lies in aggregating patient data from various CDM domains and converting them into sequences representing the patients’ medical events.
2.4 Patient Sequences and Tokenization
@@
For laboratory measurements containing continuous numerical values, we create 10 quantiles for each measurement type and assign quantile tokens ([Q1] through [Q10]) to represent the discretized values. This quantile-based approach, adapted from Ethos [7], allows us to effectively utilize numerical laboratory values within the discrete token framework.
@@ 2.5 Training
We adopt the standard Masked Language Model (MLM) method[8] to pretrain our model. To finetune the model for the downstream task of predicting chronic kidney disease prognosis, we set binary classification objectives across several follow-up periods. We define case/control groups for each task with different follow-up and assessment period combinations. We then finetune the model to predict whether a patient will progress to a specific CKD stage within the designated follow-up period. For example, one task predicts whether a patient will progress from stage 3a to stage 5 within 365 days.
@@ 3.2 Baseline Model
We compared our model against CEHR-BERT [6], a prominent and well-established benchmark within the field of electronic health record analysis. CEHR-BERT employs [masked language modeling on EHR sequences] and has demonstrated strong performance on various clinical prediction tasks. It utilizes demographic, drug, diagnosis, and procedure information.

### v2

@@ 1 Introduction
Chronic kidney disease (CKD) is a progressive condition affecting millions worldwide, up to 10% of the global population, often leading to end-stage renal failure and significant morbidity [Jadoul2024]. Accurate prognosis prediction is critical for guiding treatment strategies, optimizing resource allocation, and improving patient outcomes. Traditional prognostic models rely on statistical approaches or shallow machine learning techniques, which often struggle to capture the complex temporal and contextual relationships inherent in longitudinal clinical data.
Recent advances in deep learning, particularly transformer architectures[AttAllNeed], have revolutionized sequence modeling by leveraging self-attention mechanisms to capture long-range dependencies and contextual patterns. Unlike recurrent models like Recurrent Neural Network (RNN), transformers enable parallel processing and dynamic weighting of features, making them well-suited for heterogeneous healthcare data such as electronic health records, laboratory results, and demographic information.
Previous studies on prognosis prediction of chronic kidney disease using deep learning have primarily employed recurrent neural networks (RNNs) [Zhu2023] and transformer-based architectures [Li2020BEHRT][awan-etal-2025-predicting][pang_cehr-bert_nodate]. However, these approaches have not fully explored the integration of transformer models with laboratory values. Recently, foundational models have attempted to incorporate laboratory data, as demonstrated by Ethos [renc_zero_2024], which discretized continuous lab values into 10 quantiles for model input. Building on this idea, we adapted the decoder-based architecture of Ethos into an encoder-based model, achieving significant performance improvements in prognosis prediction.
@@ 2.3 Patient Sequences and Tokenization
For laboratory measurements containing continuous numerical values, we create 10 quantiles for each measurement type and assign quantile tokens ([Q1] through [Q10]) to represent the discretized values. This quantile-based approach, adapted from Ethos [renc_zero_2024], allows us to effectively utilize numerical laboratory values within the discrete token framework.
@@ 2.4 Training
We adopt the standard Masked Language Model (MLM) method[Devlin2019BERTPO] to pretrain our model. To finetune the model for the downstream task of predicting chronic kidney disease prognosis, we set binary classification objectives across several follow-up periods. We define case/control groups for each task with different follow-up and assessment period combinations. We then finetune the model to predict whether a patient will progress to a specific CKD stage within the designated follow-up period. For example, one task predicts whether a patient will progress from stage 3a to stage 5 within 365 days.
@@ 3.2 Baseline Model
We compared our model against CEHR-BERT [pang_cehr-bert_nodate], a prominent and well-established benchmark within the field of electronic health record analysis. CEHR-BERT employs [masked language modeling on EHR sequences] and has demonstrated strong performance on various clinical prediction tasks. It utilizes demographic, drug, diagnosis, and procedure information.